OpenStudy (jade085):

Is this process for finding derivative of R(t)= A*sqrt(t-T)/t correct? Work is already done. I just need someone to look over.

OpenStudy (jade085):

This is the work

OpenStudy (jade085):

@Preetha ?

OpenStudy (mww):

you shouldn't have squared the 2. in the second line, only the sqrt(t-T) is squared

OpenStudy (mww):

However you still arrived at the correct answer, just be careful of using parentheses...

OpenStudy (jade085):

@mww thanks

OpenStudy (kainui):

Just a tip, I like something called implicit differentiation since it makes my life easier and maybe it will for you too. I basically just rearrange it so there are no fractions are square roots with algebra from: \[R = \frac{A\sqrt{t-T}}{t}\] into \[R^2 t^2 = A^2 (t-T)\] Then I take the derivative, \[2RR't^2 + R^22t = A^2\] That whole left part I didn't even write on my paper cause I know \(R'=0\) so everything multiplying it is 0. \[R^22t=A^2\] At this point, I see I need \(R^2\) so I square the original function and plug it in, here's what \(R^2\) looks like: \[R^2 = \frac{A^2 (t-T)}{t^2}\] Now pluggin in: \[\frac{A^2 (t-T)}{t^2}2t=A^2\]\[2(t-T)=t\]\[t=2T\] I find this much less painful cause I hate fractions haha.